### v2

@@ 4.1 Implications of the Aubin property
In this subsection, we introduce a reduction approach to the Aubin property of ${\mathbb{S}}_{\rm GE}$ (1.8) characterized by the Mordukhovich criterion, for recasting it into a more accessible form. For convenience, we make the following assumption.
Assumption 4.1.
The point ${\boldsymbol{x}}^{*}$ is a nondegenerate locally optimal solution (i.e., (1.7) holds at ${\boldsymbol{x}}^{*}$) to (1.1) with ${\boldsymbol{y}}^{*}=({\bm{\lambda}}^{*},{\bm{\mu}}^{*})$ being the associated multiplier, and the index set $\{1,\ldots,J\}$ are classified by (4.1).
Thanks to Lemma 2.5, we can get the following coderivative-based characterization for the Aubin property of ${{\mathbb{S}}}_{\rm GE}$ (1.8).
@@
Under Assumption 4.1, the mapping ${{\mathbb{S}}}_{\rm GE}$ in (1.8) has the Aubin property at the origin for ${\boldsymbol{x}}^{*}$ if and only if for all ${\boldsymbol{d}}\in{\rm aff}({\mathcal{C}}({\boldsymbol{x}}^{*}))\backslash\{{\boldsymbol{0}}\}$ it holds that
$\begin{array}[]{lr}\displaystyle\nabla_{{\boldsymbol{x}}{\boldsymbol{x}}}^{2}{\mathcal{L}}({\boldsymbol{x}}^{*},{\bm{\lambda}}^{*},{\bm{\mu}}^{*}){\boldsymbol{d}}\notin{\sf rge}\nabla h({\boldsymbol{x}}^{*})\\[2.84526pt] \qquad\qquad\displaystyle-\left\{\sum_{j=1}^{J}\nabla g^{j}({\boldsymbol{x}}^{*}){\boldsymbol{p}}^{j}\ \Big{|}\ {\boldsymbol{p}}^{j}\in D^{*}{\mathcal{N}}_{{\mathcal{Q}}_{j}}(g^{j}({\boldsymbol{x}}^{*}),-({\bm{\mu}}^{*})^{j})({\mathcal{J}}g^{j}({\boldsymbol{x}}^{*}){\boldsymbol{d}})\right\}.\end{array}$ (4.8)
Proof.
Since Assumption 4.1 holds, according to Lemmas 2.4 and 2.5 we know that the mapping ${\mathbb{S}}_{\rm GE}$ in (1.8) has the Aubin property at the origin for ${\boldsymbol{x}}^{*}$ if and only if for all ${\boldsymbol{d}}\,\in{\mathds{R}}^{n}\setminus\{{\boldsymbol{0}}\}$,
$\begin{array}[]{lll}\nabla_{{\boldsymbol{x}}{\boldsymbol{x}}}^{2}{\mathcal{L}}({\boldsymbol{x}}^{*},{\bm{\lambda}}^{*},{\bm{\mu}}^{*}){\boldsymbol{d}}&\notin&-\nabla h({\boldsymbol{x}}^{*}){\mathcal{D}}^{*}{\mathcal{N}}_{\{{\boldsymbol{0}}\}}(h({\boldsymbol{x}}^{*}),-{\bm{\lambda}}^{*})({\mathcal{J}}h({\boldsymbol{x}}^{*}){\boldsymbol{d}})\\ &&-\sum_{j=1}^{J}\nabla g^{j}({\boldsymbol{x}}^{*}){\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}_{j}}(g^{j}({\boldsymbol{x}}^{*}),-({\bm{\mu}}^{*})^{j})({\mathcal{J}}g^{j}({\boldsymbol{x}}^{*}){\boldsymbol{d}}).\end{array}$ (4.9)
From (2.4) we know that $D^{*}{\mathcal{N}}_{\{{\boldsymbol{0}}\}}(h({\boldsymbol{x}}^{*}),-{\bm{\lambda}}^{*})({\mathcal{J}}h({\boldsymbol{x}}^{*}){\boldsymbol{d}})={\mathds{R}}^{m}$, if ${\mathcal{J}}h({\boldsymbol{x}}^{*}){\boldsymbol{d}}={\boldsymbol{0}}$, or else it is an empty set. Therefore, (4.9) holds whenever ${\mathcal{J}}h({\boldsymbol{x}}^{*}){\boldsymbol{d}}\neq{\boldsymbol{0}}$. Next, if ${\boldsymbol{0}}\neq{\boldsymbol{d}}\notin{\rm aff}({\mathcal{C}}({\boldsymbol{x}}^{*}))$ satisfies ${\mathcal{J}}h({\boldsymbol{x}}^{*}){\boldsymbol{d}}={\boldsymbol{0}}$, according to (4.4), one can always choose an index $j$ such that at least one of the following three conditions holds:
@@
If $j\in J_{1}$ and ${\mathcal{J}}g^{j}({\boldsymbol{x}}^{*}){\boldsymbol{d}}\neq{\boldsymbol{0}}$, by (4.1) one has $g^{j}({\boldsymbol{x}}^{*})={\boldsymbol{0}}$ and $({\bm{\mu}}^{*})^{j}\in{\sf int}{\mathcal{Q}}_{j}$. Then one can get ${\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}_{j}}(g^{j}({\boldsymbol{x}}^{*}),-({\bm{\mu}}^{*})^{j})({\mathcal{J}}g^{j}({\boldsymbol{x}}^{*}){\boldsymbol{d}})=\emptyset$ by Part (1) of Lemma 3.2 (or (2.3) if $r_{j}=0$).
If $j\in J_{4}$ and ${\mathcal{J}}g^{j}({\boldsymbol{x}}^{*}){\boldsymbol{d}}\notin{\mathds{R}}(({\dot{\mu}^{*}})^{j};-(\bar{{\bm{\mu}}}^{*})^{j})$ holds, one can get from (4.1) that $g^{j}({\boldsymbol{x}}^{*})={\boldsymbol{0}}$ and $({\bm{\mu}}^{*})^{j}\in{\sf bdry}{\mathcal{Q}}_{j}\setminus\{{\boldsymbol{0}}\}$, so that $(\dot{\mu}^{*})^{j}=\|(\bar{\bm{\mu}}^{*})^{j}\|$. Then by Part $(5)$ of Lemma 3.2 one has ${\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}_{j}}(g^{j}({\boldsymbol{x}}^{*}),-({\bm{\mu}}^{*})^{j})({\mathcal{J}}g^{j}({\boldsymbol{x}}^{*}){\boldsymbol{d}})=\emptyset$.
If $j\in J_{3}$ and $\langle{\mathcal{J}}g^{j}({\boldsymbol{x}}^{*}){\boldsymbol{d}},({\bm{\mu}}^{*})^{j}\rangle\neq 0$, one has from (4.1) that $({\bm{\mu}}^{*})^{j},g^{j}({\boldsymbol{x}}^{*})\in{\sf bdry}{\mathcal{Q}}_{j}\setminus\{{\boldsymbol{0}}\}$. Then one has ${\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}_{j}}(g^{j}({\boldsymbol{x}}^{*}),-({\bm{\mu}}^{*})^{j})({\mathcal{J}}g^{j}({\boldsymbol{x}}^{*}){\boldsymbol{d}})=\emptyset$ by Part (3) of Lemma 3.2.
In summary, according to the above discussions, we know that (4.9) holds automatically whenever ${\boldsymbol{d}}\notin{\rm aff}({\mathcal{C}}({\boldsymbol{x}}^{*}))\backslash\{{\boldsymbol{0}}\}$. Therefore, (4.8) is equivalent to (4.9). ∎
The following lemma, as an application of the reduction approach developed in [5, Section 3.4.4] for constraint systems, provides a convenient tool for characterizing the affine hull of the critical cone (4.4) via the null space of a given matrix.
@@
$\Xi:=\left({\mathcal{J}}h({\boldsymbol{x}}^{*});\,\{{\mathcal{J}}g^{j}({\boldsymbol{x}}^{*})\}_{j\in J_{1}};\,\{[({\bm{\mu}}^{*})^{j}]^{\top}{\mathcal{J}}g^{j}({\boldsymbol{x}}^{*})\}_{j\in J_{3}};\,\{S_{j}^{\top}{\mathcal{J}}g^{j}({\boldsymbol{x}}^{*})\}_{j\in J_{4}}\right),$ (4.10)
where the columns of each matrix $S_{j}\in{\mathds{R}}^{(1+r_{j})\times r_{j}}$ span a basis of $\{((\dot{\mu}^{*})^{j};-(\bar{\bm{\mu}}^{*})^{j})\}^{\perp}$. Then, one has ${\rm aff}({\mathcal{C}}({\boldsymbol{x}}^{*}))=\ker\Xi$. Moreover, $\Xi$ has full row rank.
Proof.
The first conclusion comes immediately from (4.4) and (4.10). Next, we show that $\Xi$ is surjective. Let ${\bm{\zeta}}^{0}\in{\mathds{R}}^{m}$, $\{{\bm{\zeta}}^{j}\in{\mathds{R}}^{1+r_{j}}\}_{j\in J_{1}}$, $\{\zeta^{j}\in{\mathds{R}}\}_{j\in J_{3}}$, and $\{\hat{\bm{\zeta}}^{j}\in{\mathds{R}}^{r_{j}}\}_{j\in J_{4}}$ be arbitrarily given. Define
@@
One can see from (1.7) that there exist vectors ${\boldsymbol{d}}\in{\mathds{R}}^{n}$ and ${\boldsymbol{v}}^{j}\in{\sf lin}\left({\mathcal{T}}_{{\mathcal{Q}}_{j}}\big{(}g^{j}({\boldsymbol{x}}^{*})\big{)}\right)$ such that ${\mathcal{J}}h({\boldsymbol{x}}^{*}){\boldsymbol{d}}={\bm{\zeta}}^{0}$ and ${\mathcal{J}}g^{j}({\boldsymbol{x}}^{*}){\boldsymbol{d}}+{\boldsymbol{v}}^{j}={\bm{\zeta}}^{j}$ for all $j=1,\ldots,J$, where by (4.2) the linearity space of the tangent cone takes the following form
${\sf lin}({\mathcal{T}}_{{\mathcal{Q}}_{j}}(g^{j}({\boldsymbol{x}}^{*})))=\begin{cases}{\mathds{R}}^{1+r_{j}},&\mbox{ if }j\in J_{2},\\ \{{\boldsymbol{0}}\},&\mbox{ if }j\in J_{1,4,6},\\ \{{\boldsymbol{v}}\in{\mathds{R}}^{1+r_{j}}\mid\langle\bar{\boldsymbol{v}},\bar{g}^{j}({\boldsymbol{x}}^{*})\rangle-\dot{v}\dot{g}^{j}({\boldsymbol{x}}^{*})=0\},&\mbox{ if }j\in J_{3,5}.\end{cases}$ (4.11)
Then, by using (4.11) one has ${\mathcal{J}}h({\boldsymbol{x}}^{*}){\boldsymbol{d}}={\bm{\zeta}}^{0}$, $\{{\mathcal{J}}g^{j}({\boldsymbol{x}}^{*}){\boldsymbol{d}}={\bm{\zeta}}^{j}\}_{j\in J_{1}}$, $\{{\mathcal{J}}g^{j}({\boldsymbol{x}}^{*}){\boldsymbol{d}}+{\boldsymbol{v}}^{j}={\bm{\zeta}}^{j}\}_{j\in J_{3}}$ and $\{{\mathcal{J}}g^{j}({\boldsymbol{x}}^{*}){\boldsymbol{d}}={\bm{\zeta}}^{j}\}_{j\in J_{4}}$. Moreover, for all $j\in J_{3}$, it holds that
@@
Meanwhile, for all $j\in J_{4}$, it holds that $S_{j}^{\top}{\mathcal{J}}g^{j}({\boldsymbol{x}}^{*}){\boldsymbol{d}}=S_{j}^{\top}{\bm{\zeta}}^{j}=\hat{\bm{\zeta}}^{j}$. Therefore, we know that $\Xi$ has full row rank since it is surjective. This completes the proof. ∎
@@
$\Theta:=\nabla_{{\boldsymbol{x}}{\boldsymbol{x}}}^{2}{\mathcal{L}}({\boldsymbol{x}}^{*},{\bm{\lambda}}^{*},{\bm{\mu}}^{*})-\sum_{j\in J_{3}}\frac{(\dot{\mu}^{*})^{j}}{\dot{g}^{j}({\boldsymbol{x}}^{*})}\nabla g^{j}({\boldsymbol{x}}^{*})\begin{pmatrix}1&{\boldsymbol{0}}^{\top}\\ {\boldsymbol{0}}&-I_{r_{j}}\end{pmatrix}{\mathcal{J}}g^{j}({\boldsymbol{x}}^{*}).$ (4.12)
Based on Lemma 4.2, we further reduce the condition in Lemma 4.1 for characterizing the Aubin property of the solution mapping ${{\mathbb{S}}}_{\rm GE}$ (1.8).
@@
Suppose that Assumption 4.1 holds. Let $\Theta\in{\mathds{R}}^{n\times n}$ be the matrix defined in (4.12), and $H\in{\mathds{R}}^{n\times\ell}$ be a matrix with full column rank such that ${\sf rge}H={\rm aff}({\mathcal{C}}({\boldsymbol{x}}^{*}))$. Then, the mapping ${{\mathbb{S}}}_{\rm GE}$ in (1.8) has the Aubin property at the origin for ${\boldsymbol{x}}^{*}$ if and only if for all ${\bm{\nu}}\in{\mathds{R}}^{\ell}\setminus\{{\boldsymbol{0}}\}$,
$H^{\top}\Theta H{\bm{\nu}}\notin-H^{\top}\left\{\sum_{j\in J_{4,5,6}}\nabla g^{j}({\boldsymbol{x}}^{*}){\boldsymbol{p}}^{j}\Big{|}\ {\boldsymbol{p}}^{j}\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}_{j}}(g^{j}({\boldsymbol{x}}^{*}),-({\bm{\mu}}^{*})^{j})({\mathcal{J}}g^{j}({\boldsymbol{x}}^{*})H{\bm{\nu}})\right\}.$ (4.13)
Proof.
In this setting, we know from Lemma 4.1 that the mapping ${{\mathbb{S}}}_{\rm GE}$ (1.8) has the Aubin property at the origin for ${\boldsymbol{x}}^{*}$ if and only if (4.8) holds. For a given vector ${\boldsymbol{d}}\in{\rm aff}({\mathcal{C}}({\boldsymbol{x}}^{*}))\backslash\{{\boldsymbol{0}}\}$, one can get from (4.4) that ${\mathcal{J}}g^{j}({\boldsymbol{x}}^{*}){\boldsymbol{d}}={\boldsymbol{0}}$ for all $j\in J_{1}$. Then by Part (1) of Lemma 3.2 (or (2.3) if $r_{j}=0$), $D^{*}{\mathcal{N}}_{{\mathcal{Q}}_{j}}(g^{j}({\boldsymbol{x}}^{*}),-({\bm{\mu}}^{*})^{j})({\mathcal{J}}g^{j}({\boldsymbol{x}}^{*}){\boldsymbol{d}})={\mathds{R}}^{1+r_{j}}$ holds for all $j\in J_{1}$. Meanwhile, one has that $D^{*}{\mathcal{N}}_{{\mathcal{Q}}_{j}}(g^{j}({\boldsymbol{x}}^{*}),-({\bm{\mu}}^{*})^{j})({\mathcal{J}}g^{j}({\boldsymbol{x}}^{*}){\boldsymbol{d}})=\{{\boldsymbol{0}}\}$ holds for all $j\in J_{2}$ by Part (2) of Lemma 3.2 (or (2.3) if $r_{j}=0$). Furthermore, if $j\in J_{3}$, one has $\langle{\mathcal{J}}g^{j}({\boldsymbol{x}}^{*}){\boldsymbol{d}},({\bm{\mu}}^{*})^{j}\rangle=0$. Then by Part (3) of Lemma 3.2 it holds that
@@
Therefore, by the above discussions and (4.12), the condition (4.8) can be equivalently written as for all ${\boldsymbol{d}}\in{\rm aff}({\mathcal{C}}({\boldsymbol{x}}^{*}))\setminus\{{\boldsymbol{0}}\}$ it holds that
$\begin{array}[]{ll}\Theta{\boldsymbol{d}}\notin{\mathcal{R}}-\left\{\sum\limits_{j\in J_{4,5,6}}\nabla g^{j}({\boldsymbol{x}}^{*}){\boldsymbol{p}}^{j}\ \Big{|}\ {\boldsymbol{p}}^{j}\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}_{j}}(g^{j}({\boldsymbol{x}}^{*}),-({\bm{\mu}}^{*})^{j})({\mathcal{J}}g^{j}({\boldsymbol{x}}^{*}){\boldsymbol{d}})\right\},\end{array}$ (4.14)

### v3

@@ 4.1 Implications of the Aubin property
In this subsection, we introduce a reduction approach to the Aubin property of ${\mathbb{S}}_{\rm GE}$ (8) characterized by the Mordukhovich criterion, for recasting it into a more accessible form. For convenience, we make the following assumption. {assumption} The point ${\boldsymbol{x}}^{*}$ is a nondegenerate locally optimal solution (i.e., (7) holds at ${\boldsymbol{x}}^{*}$) to (1) with ${\boldsymbol{y}}^{*}=({\bm{\lambda}}^{*},{\bm{\mu}}^{*})$ being the associated multiplier, and the index set $\{1,\ldots,J\}$ are classified by (27).
Thanks to Lemma 2.6, we can get the following coderivative-based characterization for the Aubin property of ${{\mathbb{S}}}_{\rm GE}$ (8).
@@
Under Assumption 4.1, the mapping ${{\mathbb{S}}}_{\rm GE}$ in (8) has the Aubin property at the origin for ${\boldsymbol{x}}^{*}$ if and only if for all ${\boldsymbol{d}}\in{\sf aff}({\mathcal{C}}({\boldsymbol{x}}^{*}))\backslash\{{\boldsymbol{0}}\}$ it holds that
(34) $\begin{array}[]{lr}\displaystyle\nabla_{{\boldsymbol{x}}{\boldsymbol{x}}}^{2}{\mathcal{L}}({\boldsymbol{x}}^{*},{\bm{\lambda}}^{*},{\bm{\mu}}^{*}){\boldsymbol{d}}\notin{\sf rge}\nabla h({\boldsymbol{x}}^{*})\\[2.84526pt] \qquad\qquad\displaystyle-\left\{\sum_{j=1}^{J}\nabla g^{j}({\boldsymbol{x}}^{*}){\boldsymbol{p}}^{j}\ \Big{|}\ {\boldsymbol{p}}^{j}\in D^{*}{\mathcal{N}}_{{\mathcal{Q}}_{j}}(g^{j}({\boldsymbol{x}}^{*}),-({\bm{\mu}}^{*})^{j})({\mathcal{J}}g^{j}({\boldsymbol{x}}^{*}){\boldsymbol{d}})\right\}.\end{array}$
Proof 4.4.
Since Assumption 4.1 holds, according to Lemmas 2.5 and 2.6 we know that the mapping ${\mathbb{S}}_{\rm GE}$ in (8) has the Aubin property at the origin for ${\boldsymbol{x}}^{*}$ if and only if for all ${\boldsymbol{d}}\,\in{\mathds{R}}^{n}\setminus\{{\boldsymbol{0}}\}$,
(35) $\begin{array}[]{lll}\nabla_{{\boldsymbol{x}}{\boldsymbol{x}}}^{2}{\mathcal{L}}({\boldsymbol{x}}^{*},{\bm{\lambda}}^{*},{\bm{\mu}}^{*}){\boldsymbol{d}}&\notin&-\nabla h({\boldsymbol{x}}^{*}){\mathcal{D}}^{*}{\mathcal{N}}_{\{{\boldsymbol{0}}\}}(h({\boldsymbol{x}}^{*}),-{\bm{\lambda}}^{*})({\mathcal{J}}h({\boldsymbol{x}}^{*}){\boldsymbol{d}})\\ &&-\sum_{j=1}^{J}\nabla g^{j}({\boldsymbol{x}}^{*}){\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}_{j}}(g^{j}({\boldsymbol{x}}^{*}),-({\bm{\mu}}^{*})^{j})({\mathcal{J}}g^{j}({\boldsymbol{x}}^{*}){\boldsymbol{d}}).\end{array}$
From (12) we know that $D^{*}{\mathcal{N}}_{\{{\boldsymbol{0}}\}}(h({\boldsymbol{x}}^{*}),-{\bm{\lambda}}^{*})({\mathcal{J}}h({\boldsymbol{x}}^{*}){\boldsymbol{d}})={\mathds{R}}^{m}$, if ${\mathcal{J}}h({\boldsymbol{x}}^{*}){\boldsymbol{d}}={\boldsymbol{0}}$, or else it is an empty set. Therefore, (35) holds whenever ${\mathcal{J}}h({\boldsymbol{x}}^{*}){\boldsymbol{d}}\neq{\boldsymbol{0}}$. Next, if ${\boldsymbol{0}}\neq{\boldsymbol{d}}\notin{\sf aff}({\mathcal{C}}({\boldsymbol{x}}^{*}))$ satisfies ${\mathcal{J}}h({\boldsymbol{x}}^{*}){\boldsymbol{d}}={\boldsymbol{0}}$, according to (30), one can always choose an index $j$ such that at least one of the following three conditions holds:
@@
If $j\in J_{1}$ and ${\mathcal{J}}g^{j}({\boldsymbol{x}}^{*}){\boldsymbol{d}}\neq{\boldsymbol{0}}$, by (27) one has $g^{j}({\boldsymbol{x}}^{*})={\boldsymbol{0}}$ and $({\bm{\mu}}^{*})^{j}\in{\sf int}{\mathcal{Q}}_{j}$. Then one can get ${\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}_{j}}(g^{j}({\boldsymbol{x}}^{*}),-({\bm{\mu}}^{*})^{j})({\mathcal{J}}g^{j}({\boldsymbol{x}}^{*}){\boldsymbol{d}})=\emptyset$ by Part (1) of Lemma 3.2 (or (11) if $r_{j}=0$).
If $j\in J_{4}$ and ${\mathcal{J}}g^{j}({\boldsymbol{x}}^{*}){\boldsymbol{d}}\notin{\mathds{R}}(({\dot{\mu}^{*}})^{j};-(\bar{{\bm{\mu}}}^{*})^{j})$ holds, one can get from (27) that $g^{j}({\boldsymbol{x}}^{*})={\boldsymbol{0}}$ and $({\bm{\mu}}^{*})^{j}\in{\sf bdry}{\mathcal{Q}}_{j}\setminus\{{\boldsymbol{0}}\}$, so that $(\dot{\mu}^{*})^{j}=\|(\bar{\bm{\mu}}^{*})^{j}\|$. Then by Part $(5)$ of Lemma 3.2 one has ${\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}_{j}}(g^{j}({\boldsymbol{x}}^{*}),-({\bm{\mu}}^{*})^{j})({\mathcal{J}}g^{j}({\boldsymbol{x}}^{*}){\boldsymbol{d}})=\emptyset$.
If $j\in J_{3}$ and $\langle{\mathcal{J}}g^{j}({\boldsymbol{x}}^{*}){\boldsymbol{d}},({\bm{\mu}}^{*})^{j}\rangle\neq 0$, one has from (27) that $({\bm{\mu}}^{*})^{j},g^{j}({\boldsymbol{x}}^{*})\in{\sf bdry}{\mathcal{Q}}_{j}\setminus\{{\boldsymbol{0}}\}$. Then one has ${\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}_{j}}(g^{j}({\boldsymbol{x}}^{*}),-({\bm{\mu}}^{*})^{j})({\mathcal{J}}g^{j}({\boldsymbol{x}}^{*}){\boldsymbol{d}})=\emptyset$ by Part (3) of Lemma 3.2.
In summary, according to the above discussions, we know that (35) holds automatically whenever ${\boldsymbol{d}}\notin{\sf aff}({\mathcal{C}}({\boldsymbol{x}}^{*}))\backslash\{{\boldsymbol{0}}\}$. Therefore, (34) is equivalent to (35).
The following lemma, as an application of the reduction approach developed in [5, Section 3.4.4] for constraint systems, provides a convenient tool for characterizing the affine hull of the critical cone (30) via the null space of a given matrix.
@@
(36) $\Xi:=\left({\mathcal{J}}h({\boldsymbol{x}}^{*});\,\{{\mathcal{J}}g^{j}({\boldsymbol{x}}^{*})\}_{j\in J_{1}};\,\{[({\bm{\mu}}^{*})^{j}]^{\top}{\mathcal{J}}g^{j}({\boldsymbol{x}}^{*})\}_{j\in J_{3}};\,\{S_{j}^{\top}{\mathcal{J}}g^{j}({\boldsymbol{x}}^{*})\}_{j\in J_{4}}\right),$
where the columns of each matrix $S_{j}\in{\mathds{R}}^{(1+r_{j})\times r_{j}}$ span a basis of $\{((\dot{\mu}^{*})^{j};-(\bar{\bm{\mu}}^{*})^{j})\}^{\perp}$. Then, one has ${\sf aff}({\mathcal{C}}({\boldsymbol{x}}^{*}))={\sf ker}\Xi$. Moreover, $\Xi$ has full row rank.
Proof 4.6.
The first conclusion comes immediately from (30) and (36). Next, we show that $\Xi$ is surjective. Let ${\bm{\zeta}}^{0}\in{\mathds{R}}^{m}$, $\{{\bm{\zeta}}^{j}\in{\mathds{R}}^{1+r_{j}}\}_{j\in J_{1}}$, $\{\zeta^{j}\in{\mathds{R}}\}_{j\in J_{3}}$, and $\{\hat{\bm{\zeta}}^{j}\in{\mathds{R}}^{r_{j}}\}_{j\in J_{4}}$ be arbitrarily given. Define
@@
One can see from (7) that there exist vectors ${\boldsymbol{d}}\in{\mathds{R}}^{n}$ and ${\boldsymbol{v}}^{j}\in{\sf lin}\left({\mathcal{T}}_{{\mathcal{Q}}_{j}}\big{(}g^{j}({\boldsymbol{x}}^{*})\big{)}\right)$ such that ${\mathcal{J}}h({\boldsymbol{x}}^{*}){\boldsymbol{d}}={\bm{\zeta}}^{0}$ and ${\mathcal{J}}g^{j}({\boldsymbol{x}}^{*}){\boldsymbol{d}}+{\boldsymbol{v}}^{j}={\bm{\zeta}}^{j}$ for all $j=1,\ldots,J$, where by (28) the linearity space of the tangent cone takes the following form
(37) ${\sf lin}({\mathcal{T}}_{{\mathcal{Q}}_{j}}(g^{j}({\boldsymbol{x}}^{*})))=\begin{cases}{\mathds{R}}^{1+r_{j}},&\mbox{ if }j\in J_{2},\\ \{{\boldsymbol{0}}\},&\mbox{ if }j\in J_{1,4,6},\\ \{{\boldsymbol{v}}\in{\mathds{R}}^{1+r_{j}}\mid\langle\bar{\boldsymbol{v}},\bar{g}^{j}({\boldsymbol{x}}^{*})\rangle-\dot{v}\dot{g}^{j}({\boldsymbol{x}}^{*})=0\},&\mbox{ if }j\in J_{3,5}.\end{cases}$
Then, by using (37) one has ${\mathcal{J}}h({\boldsymbol{x}}^{*}){\boldsymbol{d}}={\bm{\zeta}}^{0}$, $\{{\mathcal{J}}g^{j}({\boldsymbol{x}}^{*}){\boldsymbol{d}}={\bm{\zeta}}^{j}\}_{j\in J_{1}}$, $\{{\mathcal{J}}g^{j}({\boldsymbol{x}}^{*}){\boldsymbol{d}}+{\boldsymbol{v}}^{j}={\bm{\zeta}}^{j}\}_{j\in J_{3}}$ and $\{{\mathcal{J}}g^{j}({\boldsymbol{x}}^{*}){\boldsymbol{d}}={\bm{\zeta}}^{j}\}_{j\in J_{4}}$. Moreover, for all $j\in J_{3}$, it holds that
@@
Meanwhile, for all $j\in J_{4}$, it holds that $S_{j}^{\top}{\mathcal{J}}g^{j}({\boldsymbol{x}}^{*}){\boldsymbol{d}}=S_{j}^{\top}{\bm{\zeta}}^{j}=\hat{\bm{\zeta}}^{j}$. Therefore, we know that $\Xi$ has full row rank since it is surjective. This completes the proof.
@@
(38) $\Theta:=\nabla_{{\boldsymbol{x}}{\boldsymbol{x}}}^{2}{\mathcal{L}}({\boldsymbol{x}}^{*},{\bm{\lambda}}^{*},{\bm{\mu}}^{*})-\sum_{j\in J_{3}}\frac{(\dot{\mu}^{*})^{j}}{\dot{g}^{j}({\boldsymbol{x}}^{*})}\nabla g^{j}({\boldsymbol{x}}^{*})\begin{pmatrix}1&{\boldsymbol{0}}^{\top}\\ {\boldsymbol{0}}&-I_{r_{j}}\end{pmatrix}{\mathcal{J}}g^{j}({\boldsymbol{x}}^{*}).$
Based on Lemma 4.5, we further reduce the condition in Lemma 4.3 for characterizing the Aubin property of the solution mapping ${{\mathbb{S}}}_{\rm GE}$ (8).
@@
Suppose that Assumption 4.1 holds. Let $\Theta\in{\mathds{R}}^{n\times n}$ be the matrix defined in (38), and $H\in{\mathds{R}}^{n\times\ell}$ be a matrix with full column rank such that ${\sf rge}H={\sf aff}({\mathcal{C}}({\boldsymbol{x}}^{*}))$. Then, the mapping ${{\mathbb{S}}}_{\rm GE}$ in (8) has the Aubin property at the origin for ${\boldsymbol{x}}^{*}$ if and only if for all ${\bm{\nu}}\in{\mathds{R}}^{\ell}\setminus\{{\boldsymbol{0}}\}$,
(39) $H^{\top}\Theta H{\bm{\nu}}\notin-H^{\top}\left\{\sum_{j\in J_{4,5,6}}\nabla g^{j}({\boldsymbol{x}}^{*}){\boldsymbol{p}}^{j}\Big{|}\ {\boldsymbol{p}}^{j}\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}_{j}}(g^{j}({\boldsymbol{x}}^{*}),-({\bm{\mu}}^{*})^{j})({\mathcal{J}}g^{j}({\boldsymbol{x}}^{*})H{\bm{\nu}})\right\}.$
Proof 4.8.
In this setting, we know from Lemma 4.3 that the mapping ${{\mathbb{S}}}_{\rm GE}$ (8) has the Aubin property at the origin for ${\boldsymbol{x}}^{*}$ if and only if (34) holds. For a given vector ${\boldsymbol{d}}\in{\sf aff}({\mathcal{C}}({\boldsymbol{x}}^{*}))\backslash\{{\boldsymbol{0}}\}$, one can get from (30) that ${\mathcal{J}}g^{j}({\boldsymbol{x}}^{*}){\boldsymbol{d}}={\boldsymbol{0}}$ for all $j\in J_{1}$. Then by Part (1) of Lemma 3.2 (or (11) if $r_{j}=0$), $D^{*}{\mathcal{N}}_{{\mathcal{Q}}_{j}}(g^{j}({\boldsymbol{x}}^{*}),-({\bm{\mu}}^{*})^{j})({\mathcal{J}}g^{j}({\boldsymbol{x}}^{*}){\boldsymbol{d}})={\mathds{R}}^{1+r_{j}}$ holds for all $j\in J_{1}$. Meanwhile, one has that $D^{*}{\mathcal{N}}_{{\mathcal{Q}}_{j}}(g^{j}({\boldsymbol{x}}^{*}),-({\bm{\mu}}^{*})^{j})({\mathcal{J}}g^{j}({\boldsymbol{x}}^{*}){\boldsymbol{d}})=\{{\boldsymbol{0}}\}$ holds for all $j\in J_{2}$ by Part (2) of Lemma 3.2 (or (11) if $r_{j}=0$). Furthermore, if $j\in J_{3}$, one has $\langle{\mathcal{J}}g^{j}({\boldsymbol{x}}^{*}){\boldsymbol{d}},({\bm{\mu}}^{*})^{j}\rangle=0$. Then by Part (3) of Lemma 3.2 it holds that
@@
Therefore, by the above discussions and (38), the condition (34) can be equivalently written as for all ${\boldsymbol{d}}\in{\sf aff}({\mathcal{C}}({\boldsymbol{x}}^{*}))\setminus\{{\boldsymbol{0}}\}$ it holds that
(40) $\begin{array}[]{ll}\Theta{\boldsymbol{d}}\notin{\mathcal{R}}-\left\{\sum\limits_{j\in J_{4,5,6}}\nabla g^{j}({\boldsymbol{x}}^{*}){\boldsymbol{p}}^{j}\ \Big{|}\ {\boldsymbol{p}}^{j}\in{\mathcal{D}}^{*}{\mathcal{N}}_{{\mathcal{Q}}_{j}}(g^{j}({\boldsymbol{x}}^{*}),-({\bm{\mu}}^{*})^{j})({\mathcal{J}}g^{j}({\boldsymbol{x}}^{*}){\boldsymbol{d}})\right\},\end{array}$